An investor spent a total of $2400 in two mutual funds. One earned a 6% profit while an other earned a 4% profit. if the investor's total profit was $116 how much was invested in each mutual fund?

Answer by checkley79(3341)

.06X+.04(2400-X)=116
.06X+96-.04X=116
.02X=116-96
.02X=20
X=20/.02
X=$1,000 INVESTED @ 6%
2400-1000=1400 INVESTED @ 4%.
PROOF:
.06*1000+.04*1400=116
60+56=116
116=116